OpenStudy (anonymous):

how am I supposed to find the LCM oof this one?: 1/h* [1/(x+h)^2 - 1/x^2]

OpenStudy (anonymous):

you mean how to subtract?

OpenStudy (anonymous):

No...i just have to find the LCM.

OpenStudy (anonymous):

\[\frac{1}{(x+h)^2}-\frac{1}{x^2}=\frac{x^2-(x+h)^2}{x^2(x+h)^2}\]

OpenStudy (anonymous):

lcm??????????????? :(

OpenStudy (anonymous):

i don't know what you mean by "least common multiple" of one expression

OpenStudy (anonymous):

the least common multple of your denominators \[x^2\] and \[(x+h)^2\] is there product, \[x^2(x+h)^2\]

OpenStudy (anonymous):

aha...tx pal